# Quantum phase estimation based filtering: performance analysis and application to low-energy spectral calculation

### v1

@@ Abstract
Filtering is an important technique in quantum computing used for isolating or enhancing some specific states of quantum many-body systems. In this paper, we analyze the performance of filters based on the quantum phase estimation (QPE) algorithm, in which filtering removes states associated with bitstrings in the ancilla register above a given threshold. We show that when the conventional rectangular window function is used for the QPE input state, the resulting filter function exhibits an oscillating behavior known as the Gibbs phenomenon. We also show that in the case of the sine and Kaiser windows, this phenomenon is suppressed. Furthermore, we perform numerical simulations to compare the number of necessary queries to the Hamiltonian time evolution operation of for the QPE-based filtering algorithm and the quantum eigenvalue transformation of unitary matrices with real polynomials (QETU). We find that the number of queries required for Kaiser window-based filtering is comparable to that for QETU with optimized phase angles. As an application of the QPE-based filter, we also study a two-step algorithm for low-energy spectral simulations, composed of a coarse grid for filtering and a fine grid for obtaining final high-resolution spectra. As a benchmark of the proposed scheme for realistic continuous spectra, we present the density-of-states (DOS) calculation of antiferromagnetic type-II MnO in a one-particle approximation.
@@
Calculation of the electronic structure of complex materials is one of the most promising industrial applications of quantum computing. The last decade has seen rapid development of quantum-classical hybrid approaches that leverage currently available noisy intermediate-scale quantum computers, such as the variational quantum eigensolver [1] and recently-proposed sampling-based approaches [2, 3]. In addition, with the development of quantum hardware and quantum error correction techniques, there has also been remarkable progress on the development of algorithms designed for fault-tolerant quantum computers [4], such as quantum phase estimation (QPE) [5]. These fault-tolerant quantum algorithms are expected to yield true quantum advantage, that is, exponential speedup over classical algorithms.
Many industrially and fundamentally important phenomena in chemistry and materials science occur at a small energy scale involving the electronic ground state and low-lying excited states of a material. For example, the band gaps of typical semiconductors are of the order of $1$ eV, and the energy scale of many relevant types of chemical reactions, high-temperature superconductivity, and spintronics are in meV range. On the other hand, the total energy and the L2 norm of the Hamiltonian $\|H\|_{2}$ of materials scale with the size of the system and even for small molecules exceed $10-100$ eV. As the Hilbert space of the Hamiltonian increases, the ratio of the energy region of interest to $\|H\|_{2}$ becomes smaller. As a result, it becomes increasingly difficult to obtain low-energy spectra of the system with existing quantum algorithms, as observed in recent sampling-based quantum-classical hybrid calculations [3, 6].
To overcome this fundamental problem, quantum filtering algorithms have been developed to obtain the desired quantum state from a given input state. Examples of the ground-state filtering include the cosine [7], Gaussian [8, 9], and Lorentzian [10]-based filtering schemes. Quantum signal processing (QSP) is a versatile tool to encode arbitrary polynomial functions on a quantum circuit with a set of single-qubit rotation angles determined to construct a specified polynomial function, which offers another way of filtering [11]. Quantum eigenvalue transformation of unitary matrices with real polynomials (QETU) [12] is proposed as a simple QSP-based filtering technique with only one ancilla qubit, using controlled Hamiltonian evolution as a building block.
Utilizing QPE is yet another way of filtering quantum states. In early work, Poulin and Wocjan [13] proposed a ground-state preparation scheme based on a repeated backward operation of QPE. A separate systematic approach of filtering general energy intervals has been proposed in Ref. [14]. More recently, in Ref. [15] it is reported that the QPE-based approach with a coarse frequency grid filters out the excited states of a simple model system more efficiently than QETU. The QPE-based approach requires more ancilla qubits than QETU, but has an advantage that no procedure for finding QSP angles is needed.
@@
Here we focus on the QPE-based filtering approach, combined with the window function techniques for QPE. Recently, the window function techniques have been attracting much interest in QPE. We perform a detailed analysis on how the quality of the QPE-based filter depends on the three window functions: (1) the rectangular window, (2) the sine window, and (3) the Kaiser window. The rectangular window, which is used in the conventional QPE algorithm, is known to suffer from the spectral leakage problem [16, 17]. The sine window is known to mitigate the problem and has been applied in quantum metrology [18, 19, 20, 21] to achieve the so-called Heisenberg-limit scaling. The Kaiser window [22] is proposed as an approximation of the prolate spheroidal window or the Slepian window [23, 24, 25, 26], and recently advantages of using this window over QSP-based approaches have been reported [27, 28].
Our main focus is not to increase the probability of obtaining the ground state of the system, as done in many ground-state preparation techniques, but to increase the probability of the states in the specified small energy region without altering their relative amplitudes. This requirement is important when calculating the low-energy spectral properties of the system.
We show that the rectangular window-based filter exhibits an oscillation known as the Gibbs phenomenon, and the oscillation is mitigated when the sine and Kaiser windows are used. The error in the Kaiser window-based filter decreases exponentially with the QPE frequency grid points, which leads to the same $O(\delta^{-1}\log\epsilon^{-1})$ scaling as the QETU-based filter. Here $\delta$ and $\epsilon$ are the width of the transitional region and the error parameter, respectively. We also numerically demonstrate that the Kaiser window-based filter and QETU with phase angles determined via convex optimization require a comparable number of Hamiltonian evolution operations for a given set of ($\delta,\epsilon$), indicating that the Kaiser window-based filter can be used as a potential alternative to the QETU-based filter. Finally, we consider an application of the proposed filtering approach to the QPE-based calculation of spectral functions [29, 30, 31, 32, 33, 34, 35]. Specifically, we consider a two-step QPE approach similar to the one proposed in Ref. [15], that uses a coarse QPE grid for filtering and a fine grid for obtaining high-resolution spectra. We test the present scheme by calculating the density-of-states (DOS) of MnO in a one-particle approximation [36] as an example of realistic spectra.
@@ II.1 General formalism
where $\phi_{\mu}$ are the eigenstates of the Hamiltonian $H$ with eigenenergies $E_{\mu}$, and $C_{\mu}$ are expansion coefficients normalized as $\sum_{\mu}|C_{\mu}|^{2}=1$. After the state-preparation step, the controlled time evolution $U_{HT}=\textrm{exp}[iHT]$ is performed, where $T$ is a time duration chosen to scale the energy range of the system in $[0,2\pi]$. Throughout this work $\hbar=1$. Finally, the inverse of quantum Fourier transform (QFT) is performed. The quantum state before the measurement in Fig. 1 can be represented as
@@
are normalized $\mu$-dependent states. Thus, the QPE-based filtering amounts to extracting the first term of Eq. (8). One way to do this truncation is a coherent (i.e., fully quantum) filtering without measurement, by first performing an ideal quantum amplitude amplification (QAA) [37, 38, 39] and then tracing out the ancilla register. Another approach is through classical post-selection of the ancilla states after the measurement, as depicted in Fig. 1. In this case, the result becomes probabilistic, and the success probability can be increased by performing QAA on the corresponding states. It should be noted that that the resulting outcome in this classical filtering approach is a probability distribution, rather than a quantum state (the first term of Eq. (8)). This approach therefore differs from usual quantum filters but can be applied to enhance the spectral estimation, as discussed in Section III.
@@
In both quantum and classical filtering approaches, the net effect of filtering is represented as the renormalization of the weight of each state $|\phi_{\mu}\rangle$ (see the Appendix for additional explanation on the classical approach):
@@
Although it is not an essential point of this approach, in this work we choose $y_{c}$ as $y_{c}=2^{n-m}-1$ with $m(<n)$ a positive integer. With this choice, $\omega_{y_{c}}=\frac{2\pi}{NT}y_{c}\approx\frac{2\pi}{T}\frac{1}{2^{m}}$, and we post-select or amplify the states whose first $m$ bits in the ancilla register are 00 (ex. $|{\bf{00}}1101\rangle$ when $m=2$). This filtering corresponds to a low-pass filter, but it is straightforward to consider filtering for general frequency intervals [14].
@@ II.2 Window functions
The performance of the QPE-based filtering depends crucially on the form of $a_{j}$. In this work, we consider three forms of $a_{j}$: (1) the rectangular window, (2) the sine window, and (3) the Kaiser window.
@@
The second window is the sine window originally proposed in quantum metrology [18, 19, 20, 21]:
@@
In this case, the resulting RMSE is proportional to $1/N$ [20], which is known as the Heisenberg-limited scaling. The corresponding $U_{\textrm{in}}$ can be constructed in $O(n)$ complexity [40].
The third window is the Kaiser window [22, 41, 27, 28], which is introduced as an approximation of the prolate spheroidal window [23, 24, 25, 26]. We consider the following form:
@@
where $I_{0}$ is the modified Bessel function of the order zero and $\alpha$ is a positive parameter. Note that our form is slightly different from other works [41, 27, 28], in that our definition $j$ goes from 00 to $N-1$. The Kaiser window and the prolate spheroidal window are more difficult to implement than the first two, but there have been proposals [42, 26]. In particular, the method of Ref. [42] uses quantum eigenvalue transformation with $O(n\sqrt[4]{\alpha+1}(\alpha+\log\epsilon^{-1}))$ gate complexity within error $\epsilon$.
We compare the results of all three windows in Fig. 3. In Fig. 3(a) we plot $a_{j}$ for the three windows. The Kaiser window has a Gaussian-like structure more localized than the sine window, and its spread is controlled by the parameter $\alpha$. In Figs. 3(b) and 3(c) we compare $P_{\mu}(y)$ defined in Eq. (5) for the three windows, for a specific example of $|\phi_{\mu}\rangle$ with $\frac{NT}{2\pi}E_{\mu}=10.2$ and $N=64$. The $P_{\mu}(y)$ in the rectangular window has the highest peak but has large leakage, and the sine window result shows a smaller tail probability [31]. For these two windows, the expression of $A_{\mu}(y)$ in Eq. (4) can easily be obtained; for the rectangular window
@@
The properties of probability distribution for the Kaiser window have been investigated in detail [41, 27, 28]. As can be seen in Fig. 3(c), there is a clear separation between the high-probability region called mainlobe, which is confined in $\approx 2\alpha$ QPE frequency grid points near the peak center, and the rather flat region outside the mainlobe, called sidelobes. The sum of $P_{\mu}(y)$ in the sidelobes gives the amount of the leakage $\epsilon$, which is related to $\alpha$ in the leading order by [41]
@@ II.3 QPE-based filters for the three windows
Figure 5: (a) Weight-renormalization factor (filter) $\mathcal{R}_{\mu}$ (Eq. (9)) for the rectangular and the sine windows (top) and for the Kaiser window with $\alpha=3$ and $6$ (bottom). (b) Deviations of $\mathcal{R}_{\mu}$ from the ideal filter, $\Delta\mathcal{R}_{\mu}$ given in Eq. (14), on the logarithmic scale for $E_{\mu}\leq 2$ (left) and $E_{\mu}\geq 5$ (right). The vertical dotted line in the left panel shows the position of $\omega_{y_{c}}$, and the filled squares indicate the positions of the QPE frequency grid points $\omega_{y}$.
Figure 5(a) shows $\mathcal{R}_{\mu}$ (given in Eq. (9)) as a function of $E_{\mu}$ for the three window functions described in the previous section. The parameters used are $T=1$, $n=6$ ($N=64$) and $y_{c}=15$, corresponding to extracting the states in approximately one fourth of the entire energy range. The rectangular-window-based filter exhibits an oscillating behavior known as the Gibbs phenomenon [43], which originates from the discontinuity of the truncated frequency spectrum. In the case of the other two windows this oscillation is suppressed at this scale. As shown in the insets of Figs. 5(a) and 5(b), compared to the rectangular and sine window-based filters, the Kaiser window-based filters have a larger transitional region where $\mathcal{R}_{\mu}$ deviates from one and zero. This reflects the fact that the probability distribution of the Kaiser window has a broader mainlobe as shown in Figs. 3(b) and 3(c).
In Fig. 5(b), the errors of these filters, $\Delta\mathcal{R}_{\mu}$ defined in Eq. (14), are plotted together with the QPE frequency grid points (filled squares). The sine window result also shows oscillation, though of smaller magnitude than the rectangular window. By analyzing $P_{\mu}(y)$ [31], it can be shown that these two filters are exact (i.e. $\Delta\mathcal{R}_{\mu}=0$) for $E_{\mu}=\omega_{y^{\prime}}$ and $E_{\mu}=\omega_{y^{\prime}+1/2}$ for the rectangular and sine windows, respectively, with $y^{\prime}$ any integer.
@@
This is the number of time evolution operations $U_{HT}$, required to filter the state with given error $\epsilon$.

### v3

@@ Abstract
Filtering refers to the process of using algorithms to enhance desired or mask undesired states in quantum computing, and is essential for preparing or isolating quantum many-body states. One class of filters is based on quantum phase estimation (QPE). Such QPE-based filters have a simple circuit implementation, but their performance remains unclear, particularly when compared to other filtering methods such as those with quantum signal processing (QSP). In this work, we investigate the performance of QPE-based filters with three commonly used window functions: rectangular window, sine window, and Kaiser window. We analyze how choice of window affects results; one noticeable observation is that the oscillation of the filter (the Gibbs phenomenon) seen in the conventional rectangular window result can be mitigated when the sine and Kaiser windows are applied. Furthermore, we compare the performance of QPE-based methods to quantum eigenvalue transformation of unitary matrices with real polynomials (QETU), and find that the Kaiser window-based filter and QETU result in a similar number of queries to the Hamiltonian time evolution operation. QPE-based filtering methods may therefore be an alternative to signal-processing-based methods. Finally, as an application of the QPE-based filter, we propose a two-step QPE algorithm for low-energy spectral simulations, composed of a coarse grid for filtering and a fine grid for obtaining final high-resolution spectra. As a benchmark of the proposed scheme for realistic continuous spectra, we present the density-of-states (DOS) calculation of antiferromagnetic type-II MnO in a one-particle approximation.
@@
Calculation of electronic structure of complex materials is one of the most promising industrial applications of quantum computing. The last decade has seen rapid development of quantum-classical hybrid approaches that leverage currently available noisy intermediate-scale quantum computers, such as the variational quantum eigensolver [Peruzzo2014] and sampling-based approaches [kanno2023quantumselectedconfigurationinteractionclassical, robledomoreno2024chemistryexactsolutionsquantumcentric]. As advances in fault-tolerant hardware and error correction techniques continue, fully-quantum algorithms [zhang2025faulttolerantquantumalgorithmsquantum] such as quantum phase estimation (QPE) [kitaev1995quantummeasurementsabelianstabilizer] promise true quatum advantage in the form of exponential speed up over classical computers.
However, even with fault-tolerant hardware, preparing the precise ground state and resolving low-energy spectra are crucial, as many phenomena of industrial and academic importance occur at the small energy scales between the electronic ground state and low-lying excited states. For example, semiconductor band gaps are typically on the order of $1$ eV [yu2010fundamentals], and the energy scale of many relevant types of chemical reactions, high-temperature superconductivity [brooks2007handbook], and spintronics [RevModPhys.76.323, spintronics2] are in the meV range. On the other hand, the total energy and the L2 norm of the Hamiltonian $\|H\|_{2}$ are extensive, and even for small molecules exceed 1010010-100 eV. This difference in energy scales not only makes the ground-state preparation difficult, as shown in recent sample-based calculations [robledomoreno2024chemistryexactsolutionsquantumcentric, D5CP02202A, yu2025quantumcentricalgorithmsamplebasedkrylov], but also makes it hard to resolve small low-energy excitations in quantum algorithms such as QPE.
To overcome this fundamental problem, quantum filtering algorithms have been developed to obtain the desired quantum state from a given input state. Examples of the ground-state filtering include the cosine [10.1063/1.5027484], Gaussian [PhysRevA.106.032420, Wang2022statepreparation], and Lorentzian [Irmejs2024efficientquantum]-based filtering schemes. There are also filtering methods based on quantum signal processing (QSP) [Low2019hamiltonian, PhysRevLett.118.010501], such as quantum singular value transformation (QSVT) [10.1145/3313276.3316366, PRXQuantum.2.040203, Lin2020nearoptimalground] and quantum eigenvalue transformation of unitary matrices with real polynomials (QETU) [PRXQuantum.3.040305]. However, in all QSP-based methods one has to find a set of single qubit phase angles to construct the target function on a classical computer. This procedure requires two steps. First, one needs to find a good polynomial approximation of the target function. Second, the phase factors realizing this polynomial function must be determined numerically. Although classical algorithms have been developed for the angle-finding procedure [Haah2019product, chao2020findinganglesquantumsignal, PhysRevA.103.042419, Wang2022energylandscapeof, Ying2022stablefactorization], it is worth looking for a simpler alternative filtering approach that avoids this procedure.
Here we choose to study filters based on QPE, a class of algorithms that leverages the fact that eigenenergies of the system are approximately encoded in the ancilla register in binary expansion form. In an early work, Poulin and Wocjan [PhysRevLett.102.130503] detail a ground-state preparation scheme based on a repeated backward operation of QPE, and in Ref. [Jensen_2021], P. W. K. Jense et al. propose a general systematic approach of amplifying quantum states within a specified energy interval. In addition, a recent study reported [PRXQuantum.5.040339] a simple model calculation in which their QPE-based filter with a coarse frequency grid was more cost-effective than QETU, suggesting QPE has the potential to outperform QSP-based methods. However, without a thorough understanding of the behavior and properties of QPE-based filters, it is difficult to assess any potential advantages.
We therefore aim to investigate the characteristics of the QPE-based filter through theoretical and numerical analyses, particularly the effect of the initial state of the ancilla register. This initial state, called the window function, is known to affect the performance of QPE [PhysRevA.54.4564, PhysRevLett.98.090501, 4655455, PhysRevX.8.041015, patel2024optimalcoherentquantumphase, greenaway2024casestudyqsvtassessment, PRXQuantum.6.020327], prompting questions about its role in QPE-based filters. We focus on filters that increase the probability of the states in a specified small energy region without altering their relative amplitudes, rather than increasing the probability of obtaining the ground state of the system as done in many ground-state preparation techniques. This style of filtering has various applications, such as spectroscopy simulations.
@@
We study three commonly used window functions: (1) the rectangular window, (2) the sine window, and (3) the Kaiser window. The rectangular window is the most common and used in the conventional QPE algorithm, although it is known to suffer from the spectral leakage problem [1455106, 9762511]. The sine window is known to mitigate the problem and has been applied in quantum metrology [PhysRevA.54.4564, PhysRevLett.98.090501, 4655455, PhysRevX.8.041015] to achieve the sought-after Heisenberg-limit scaling. The Kaiser window [1163349] is proposed as an approximation of the prolate spheroidal window or the Slepian window [6773659, https://doi.org/10.1002/sapm196544199, Imai_2009, patel2024optimalcoherentquantumphase], and recently advantages of using this window over QSVT-based [PRXQuantum.2.040203, greenaway2024casestudyqsvtassessment] and single-ancilla [PRXQuantum.6.020327] phase estimation algorithms have been reported.
First, we analyze the general properties of QPE-based filters. We show that the rectangular window-based filter exhibits an oscillation known as the Gibbs phenomenon, and the oscillation is mitigated when the sine and Kaiser windows are used. We also show that the effect of the window function on the filter can be incorporated into a multiplicative factor. The error in the Kaiser window-based filter decreases exponentially with the QPE frequency grid points, leading to the $O(\delta^{-1}\log\epsilon^{-1})$ query complexity. Here $\delta$ and $\epsilon$ are the width of the transitional region and the error parameter, respectively.
Next, we compare the performance of QPE-based filters and a single-ancilla QSP-based approach QETU [PRXQuantum.3.040305]. The QETU method is known to achieve nearly optimal query complexities for the ground-state energy estimation and the ground-state preparation [PRXQuantum.3.040305], and recently a demonstration on trapped-ion quantum hardware has been reported [karacan2025filterenhancedadiabaticquantumcomputing]. We find that the Kaiser window-based filter and QETU with phase angles determined via convex optimization require a comparable number of Hamiltonian evolution operations for a given set of ($\delta,\epsilon$). This result indicates that the Kaiser window-based filter can be used as a potential alternative to the QETU-based filter. We discuss possible cases where the QPE-based filter can be advantageous compared to QETU.
Finally, we consider an application of the proposed filtering approach to the QPE-based calculation of spectral functions [PhysRevC.100.034610, PhysRevA.101.012330, PhysRevA.110.022618, PhysRevA.110.L060401, fomichev2024simulatingxrayabsorptionspectroscopy, PhysRevA.92.062318, nishi2025demonstrationlogicalquantumphase]; specifically, a two-step QPE approach similar to the one proposed in Ref. [PRXQuantum.5.040339], that uses a coarse QPE grid for filtering and a fine grid for obtaining high-resolution spectra. We evaluate the performance of the proposed scheme by calculating the density-of-states (DOS) of MnO in a one-particle approximation [martin2020electronic] as an example of realistic spectra.
@@ II.1 General formalism
where $\phi_{\mu}$ are the eigenstates of the Hamiltonian $H$ with eigenenergies $E_{\mu}$, and $C_{\mu}$ are expansion coefficients normalized as $\sum_{\mu}|C_{\mu}|^{2}=1$. After the state-preparation step, the controlled time evolution $U_{HT}=\textrm{exp}[iHT]$ is performed, where $T$ is a time duration chosen to scale the energy range of the system in $0,2\pi]$. Throughout this work $\hbar=1$. Finally, the inverse of quantum Fourier transform (QFT) is performed. The quantum state before the measurement in Fig. 1 can be represented as
@@
are normalized $\mu$-dependent states. Thus, the QPE-based filtering amounts to extracting the first term of Eq. (8). One way to do this truncation is a coherent (i.e., fully quantum) filtering without measurement, by first performing an ideal quantum amplitude amplification (QAA) [gilles2002, PhysRevLett.113.210501, PRXQuantum.2.040203] and then tracing out the ancilla register. Another approach is through classical post-selection of the ancilla states after the measurement, as depicted in Fig. 1. In this case, the result becomes probabilistic, and the success probability can be increased by performing QAA on the corresponding states. This approach therefore differs from usual quantum filters but can be applied to enhance the spectral estimation, as discussed in Section III.
@@
In both quantum and classical filtering approaches, the net effect of filtering is represented as the renormalization of the weight of each state $|\phi_{\mu}\rangle$ (see the Appendix A for additional explanation on the classical approach):
@@
Although it is not an essential point of this approach, in this work we choose $y_{c}$ as $y_{c}=2^{n-m}-1$ with $m(<n)$ a positive integer. With this choice, $\omega_{y_{c}}=\frac{2\pi}{NT}y_{c}\approx\frac{2\pi}{T}\frac{1}{2^{m}}$, and we post-select or amplify the states whose first $m$ bits in the ancilla register are 0 (ex. $|{\bf{00}}1101\rangle$ when $m=2$). This filtering corresponds to a low-pass filter, but it is straightforward to consider filtering for general frequency intervals [Jensen_2021].
@@ II.2 Window functions
The performance of the QPE-based filtering depends crucially on the form of $a_{j}$. In this work, we consider three forms of $a_{j}$: (1) the rectangular window, (2) the sine window, and (3) the Kaiser window. Note that many other windows have been proposed in classical signal processing [prabhu], and a detailed exploration of their characteristics is left for future work.
@@
The second window is the sine window originally proposed in quantum metrology [PhysRevA.54.4564, PhysRevLett.98.090501, 4655455, PhysRevX.8.041015]:
@@
In this case, the resulting RMSE is proportional to $1/N$ [4655455], which is known as the Heisenberg-limited scaling. The corresponding $U_{\textrm{in}}$ can be constructed in $O(n)$ complexity [PhysRevD.106.034503].
The third window is the Kaiser window [1163349, PRXQuantum.5.010319, greenaway2024casestudyqsvtassessment, PRXQuantum.6.020327], which is introduced as an approximation of the prolate spheroidal window [6773659, https://doi.org/10.1002/sapm196544199, Imai_2009, patel2024optimalcoherentquantumphase]. We consider the following form:
@@
where $I_{0}$ is the modified Bessel function of the order zero and $\alpha$ is a positive parameter. Note that our form is slightly different from other works [PRXQuantum.5.010319, greenaway2024casestudyqsvtassessment, PRXQuantum.6.020327], in that our definition $j$ goes from 0 to N1N-1. The Kaiser window and the prolate spheroidal window are more difficult to implement than the first two, but there have been proposals [mcardle2022quantumstatepreparationcoherent, patel2024optimalcoherentquantumphase]. In particular, the method of Ref. [mcardle2022quantumstatepreparationcoherent] uses QSVT with $O(n\sqrt[4]{\pi\alpha+1}(\pi\alpha+\log\epsilon^{-1}))$ gate complexity within error $\epsilon$. In Appendix B, we review this approach and estimate the cost of constructing the Kaiser window.
We compare the results of all three windows in Fig. 3. In Fig. 3(a) we plot $a_{j}$ for the three windows. The Kaiser window has a Gaussian-like structure more localized than the sine window, and its spread is controlled by the parameter $\alpha$. In Figs. 3(b) and 3(c) we compare $P_{\mu}(y)$ defined in Eq. (5) for the three windows, for a specific example of $|\phi_{\mu}\rangle$ with $\frac{NT}{2\pi}E_{\mu}=10.2$ and $N=64$. The $P_{\mu}(y)$ in the rectangular window has the highest peak but has large leakage, and the sine window result shows a smaller tail probability [PhysRevA.110.022618]. For these two windows, the expression of $A_{\mu}(y)$ in Eq. (4) can easily be obtained; for the rectangular window
@@
The properties of probability distribution for the Kaiser window have been investigated in detail [PRXQuantum.5.010319, greenaway2024casestudyqsvtassessment, PRXQuantum.6.020327]. As can be seen in Fig. 3(c), there is a clear separation between the high-probability region called mainlobe, which is confined in $\approx 2\alpha$ QPE frequency grid points near the peak center, and the rather flat region outside the mainlobe, called sidelobes. The sum of $P_{\mu}(y)$ in the sidelobes gives the amount of the leakage $\epsilon$, which is related to $\alpha$ in the leading order by [PRXQuantum.5.010319]
@@ II.3 QPE-based filters for the three windows
Figure 5: (a) Weight-renormalization factor (filter) $\mathcal{R}_{\mu}$ (Eq. (9)) for the rectangular and the sine windows (top) and for the Kaiser window with $\alpha=3$ and 66 (bottom). (b) Deviations of $\mathcal{R}_{\mu}$ from the ideal filter, $\Delta\mathcal{R}_{\mu}$ given in Eq. (14), on the logarithmic scale for $E_{\mu}\leq 2$ (left) and $E_{\mu}\geq 5$ (right). The vertical dotted line in the left panel shows the position of $\omega_{y_{c}}$, and the filled squares indicate the positions of the QPE frequency grid points $\omega_{y}$.
Figure 5(a) shows $\mathcal{R}_{\mu}$ (given in Eq. (9)) as a function of $E_{\mu}$ for the three window functions described in the previous section. The parameters used are $T=1$, $n=6$ ($N=64$) and $y_{c}=15$, corresponding to extracting the states in approximately one fourth of the entire energy range. The rectangular-window-based filter exhibits an oscillating behavior known as the Gibbs phenomenon [Jerri1998], which originates from the discontinuity of the truncated frequency spectrum. In the case of the other two windows this oscillation is suppressed at this scale. As shown in the insets of Figs. 5(a) and 5(b), compared to the rectangular and sine window-based filters, the Kaiser window-based filters have a larger transitional region where $\mathcal{R}_{\mu}$ deviates from one and zero. This reflects the fact that the probability distribution of the Kaiser window has a broader mainlobe as shown in Figs. 3(b) and 3(c).
In Fig. 5(b), the errors of these filters, $\Delta\mathcal{R}_{\mu}$ defined in Eq. (14), are plotted together with the QPE frequency grid points (filled squares). The sine window result also shows oscillation, though of smaller magnitude than the rectangular window. By analyzing $P_{\mu}(y)$ [PhysRevA.110.022618], it can be shown that these two filters are exact (i.e. $\Delta\mathcal{R}_{\mu}=0$) for $E_{\mu}=\omega_{y^{\prime}}$ and $E_{\mu}=\omega_{y^{\prime}+1/2}$ for the rectangular and sine windows, respectively, with $y^{\prime}$ any integer.
@@
The corresponding number of ancilla qubits is $O\Bigl(\log(\delta^{-1}\log\epsilon^{-1})\Bigr)$.
In practice, from a given set of ($\epsilon,\delta,E_{\textrm{targ}}$), the parameter $\alpha$, the number of ancilla qubits $n$, and the cutoff parameter $y_{c}$ in the Kaiser window-based filter can be determined as follows: One can choose $\alpha$ from the relation between $\alpha$ and $\epsilon$ shown in Fig. 4. The width of the translational region $\delta$ and $\alpha$ are related by $\delta\approx\frac{2\pi}{2^{n}}\alpha$, from which one can determine the number of qubits $n$. The cutoff parameter $y_{c}$ can be selected from $2^{n}$ QPE grid points.